### v1

# Violations of the null convergence condition in kinematical transitions between singular and regular black holes, horizonless compact objects, and bounces

Abstract
Abstract:
How do regular black holes evade the Hawking–Penrose singularity theorems? Various models of stationary regular black holes, in particular those proposed by Bardeen, Dymnikova, Hayward, and close variants thereof, globally satisfy the null convergence condition (NCC). At first glance this might seem puzzling, as the NCC must generically be violated to avoid the focusing point implied by the singularity theorems. In fact, the extant singularity theorems depend on subtle global assumptions and do not provide information about where and when the singularity actually forms. In particular inner horizons are typically reached at finite affine parameter, before null geodesic focusing occurs, and the region inside the inner horizon is not itself a trapped region. Specifically, the Bardeen, Dymnikova, Hayward models of stationary regular black holes feature an inner Cauchy horizon which violates global hyperbolicity, hence violating one of the key assumptions of Penrose’s singularity theorem, and furthermore challenging their viability as long-living end-points of gravitational collapse. In contrast, during non-stationary processes describing kinematic transitions between standard singular black holes and regular black holes or horizonless compact objects, the inner horizon — when present — need not act as a Cauchy horizon. This raises the intriguing possibility that the NCC might instead be violated during intermediate stages of such transitions. Our detailed analysis confirms that NCC violations occur frequently during such kinematic transitions, even when the stationary end-point spacetimes respect the NCC. We also investigate analogous transitions toward black-bounce spacetimes and their horizonless compact counterparts, wormholes, where the NCC is always violated. These findings offer new insights into how regular black holes and related objects evade the constraints imposed by the standard singularity theorems.
Date: Sunday 2 February 2025 (NZT); -ed February 1, 2025.
Keywords: Null convergence; Cauchy horizons; regular black holes; black bounces
I Introduction
How do regular black holes (RBHs) evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems? This question is more subtle than one might suppose. Indeed, it is easy to check that the most common regular black holes (Bardeen, Dymnikova, Hayward, and close variants thereof Bardeen:1968bh ; Dymnikova:1992ux ; Hayward:2005gi ) globally satisfy the null convergence condition (NCC). [^1] More distantly related regular black holes, such as the Simpson–Visser black bounce Simpson:2018tsi (or the closely related Simpson–Visser naked traversable wormhole) do, however, globally violate the NCC. In contrast, the Minkowski-core regular black hole Simpson:2019mud ; Berry:2020ntz ; Simpson:2021dyo ; Simpson:2021zfl exhibits intermediate behaviour — the NCC is satisfied deep in the core and violated in the periphery. We focus on the geometric NCC, rather than the implicitly dynamical null energy condition (NEC) since the NEC makes dynamical assumptions regarding the applicability of the Einstein equations — we shall instead phrase the discussion in purely geometrical terms, eschewing the need for any particular choice of dynamics Carballo-Rubio:2019fnb . (For general discussions of energy conditions, in contrast to what can be done at a purely kinematic level, see, for instance, Curiel:2014zba ; Borde:1987qr ; Barcelo:2002bv ; Visser:1999de ; Martin-Moruno:2017exc ; Martin-Moruno:2013wfa ; Martin-Moruno:2013sfa ; Martin-Moruno:2015ena .) Since the most commonly occurring regular black holes satisfy the NCC, they must evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems in some other manner — the only remaining point of evasion being the hypothesized (global) causality condition.
Specifically, the presence of a static or stationary inner horizon, which is thereby a Cauchy horizon, violates global hyperbolicity. While improvements of the Penrose theorem, such as the Hawking–Penrose theorem, do not explicitly require the presence of a Cauchy hypersurface, still they implicitly assume some localized form of global hyperbolicity. (For instance, by postulating the absence of closed timelike curves; see e.g. Senovilla:2014gza ).
The above observations apply also to the much more familiar, and singular, Reissner–Nordström, Kerr, or Kerr–Newman spacetimes. It is perhaps a little sobering to realize that, due to the presence of Cauchy horizons, the standard Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems simply do not (directly) apply to any of the Reissner–Nordström, Kerr, or Kerr–Newman spacetimes. Certainly one can easily verify the presence of singularities in these spacetimes by direct computation, however, the presence of these singularities cannot be deduced from the standard Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems. (See in particular the closing comments in Senovilla:2014gza .)
In non-stationary spacetimes the situation is much more complicated and interesting. If, for example, one constructs a kinematic model of a transition from/to a regular black hole to/from either a standard (singular) black hole, or to/from some horizonless compact object (HCO), then the evolving inner horizon (when present) need not be a Cauchy horizon. Similarly, in a dynamical gravitational collapse leading to the creation of a Kerr black hole only an inner horizon will need to be considered at finite times, at least before the formation of a time-like singularity, (which will then also imply the formation of a Cauchy horizon).
If the inner horizon is not Cauchy, then there is no a priori reason to not apply the singularity theorems and hence the avoidance of a singularity implied by a late-time realization of one of the stationary regular black-hole geometries naturally leads to the conclusion that the NCC most probably has to be violated during the kinematic transition, though not at the stationary end-point. In this work, we shall verify that in several scenarios this is actually what happens: NCC violation can and does typically occur during the kinematic transition, even between end-point spacetimes that do themselves satisfy the NCC.
A comment is in order to better justify our specific choice of examples. Regular black holes are typically assumed to arise through some form of quantum-gravity-induced regularization that prevents the formation of a focusing point for the outgoing null congruence. [^2] Remember that in the trapped region “outgoing” means “less ingoing”. This avoidance of focusing points is a crucial step in the framework of the Hawking–Penrose singularity theorems and plays a key role in predicting the classes of regular solutions that replace the traditional singular black holes Carballo-Rubio:2019fnb ; Carballo-Rubio:2019nel ; Carballo-Rubio:2018pmi . Consequently, the most relevant transition for our study would involve a stellar-like object collapsing into one of these regular solutions.
Unfortunately, such transitions require detailed knowledge of the dynamics governing the process — knowledge that is currently unavailable. For this reason, in this work, we do not consider fully realistic collapse scenarios. Instead, we model transitions corresponding to the regularization of pre-existing spherically symmetric, static, singular solutions, such as the Schwarzschild black hole, or metrics that, while still featuring a curvature singularity, are at least continuous at $r=0$, i.e. black holes with integrable singularities, which allows us to verify that any potential artefacts, introduced by an initially ill-defined metric at the origin, do not affect the nature of the transition.
Additionally, we explore the only type of transition between a regular black hole and a horizonless object that can be easily modeled: the transition to their horizonless compact counterparts Carballo-Rubio:2022nuj . This provides a framework for analyzing how regular black holes might connect to alternative compact configurations and sheds light on the underlying mechanisms.
II Null convergence condition in spherically symmetric spacetimes
Consider a spherically symmetric spacetime with line element in ingoing Eddington–Finkelstein coordinates
[…]
$R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}\geq 0\quad\text{for {\it any} null vector}\,\,\,\ell^{\mu}.$ (2)
Let us first introduce the two radial null vectors
$k^{\mu}=-\partial_{r}=(0,-1,0,0)\,\,\,\quad\text{and}\,\,\,\quad l^{\mu}=\partial_{v}+\frac{f}{2}\partial_{r}=\quantity(1,\frac{f}{2},0,0)\,\,\,$ (3)
[…]
$R_{\mu\nu}\;k^{\mu}k^{\nu}$ $=$ $-\frac{2R^{\prime\prime}}{R}\,,$ (4)
$R_{\mu\nu}\;l^{\mu}l^{\nu}$ $=$ $-\frac{1}{2R}\quantity(2\dot{f}R^{\prime}+f^{2}R^{\prime\prime}+4f\dot{R}^{\prime}-2f^{\prime}\dot{R}+4\ddot{R})\,,$ (5)
where primes and dots denote partial derivatives with respect to $r$ and $v$. Note that the first expression corresponding to an ingoing radial null geodesic does not involve any time derivatives. The second expression, corresponding to “outgoing” radial null geodesics, is more subtle.
For completeness note that in spherical symmetry any arbitrary vector can without loss of generality be expressed in the form $\ell^{\mu}=(\alpha,\beta,\gamma,0)$, whence imposing the null condition and rescaling $\alpha\to 1$ implies
$\ell^{\mu}=\left(1,{f(r,v)-\gamma^{2}R(r,v)^{2}\over 2},\gamma,0\right)$ (6)
with $g(\ell,\ell)=0$; $g(\ell,k)=-1$; and $g(\ell,l)=-{R(r,v)^{2}\gamma^{2}\over 2}$.
Note that $\lim_{\gamma\to 0}\ell^{\mu}=l^{\mu}$ and $\lim_{\gamma\to\infty}[\gamma^{-2}\ell^{\mu}]=(0,-{1\over 2}R^{2},0,0)\propto k^{\mu}$, so this generic null vector smoothly interpolates between the two radial null vectors. The general quantity of interest in the NCC, $R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}$, is then easily computed to be:
$R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}=[R_{\mu\nu}\;l^{\mu}l^{\nu}]+{\gamma^{4}R(r,v)^{4}\over 4}[R_{\mu\nu}\;k^{\mu}k^{\nu}]+{\gamma^{2}\over 2}\left(R^{2}f^{\prime\prime}-2f\;[R^{\prime}]^{2}-4R^{\prime}\dot{R}+2\right).$ (7)
Therefore, even in the most general null directions, major insights originate from evaluating the NCC in the two radial null directions. (In the discussion below we will primarily be interested in establishing violations of the NCC, so it will often be sufficient to restrict attention to the specific quantity $R_{\mu\nu}\;l^{\mu}l^{\nu}$ corresponding to “outgoing” radial null geodesics.) Now we will distinguish two classes of geometries.
II.1 Spacetimes with $R(r,v)=r$
For $R(r,v)=r$ the expression in (4) vanishes, $R_{\mu\nu}\;k^{\mu}k^{\nu}\to 0$, whereas (5) simplifies to
$R_{\mu\nu}\;l^{\mu}l^{\nu}=-\frac{\dot{f}}{r}=\frac{2\dot{m}}{r^{2}}\,.$ (8)
We see that in any (instantaneously) static spacetime ($\dot{m}=0$) the NCC associated with the radial null vectors $k^{\mu}$ and $l^{\mu}$ is always marginally satisfied. This holds for any  static spherically symmetric black hole for which $R(r,v)=r$, i.e., in particular for the singular Schwarzschild black hole equally well as for regular black holes with smooth de Sitter cores, such as the Dymnikova Dymnikova:1992ux , Bardeen Bardeen:1968bh or Hayward Hayward:2005gi black holes. This observation motivates us to investigate kinematical models for the formation of these geometries during a transition from a singular black hole and to analyze possible necessary violations of the NCC. In any non-static situation it follows from (8) that the (radial) NCC will be violated if and only if $\dot{m}<0$.
Even in a completely general null direction $\ell^{\mu}=\left(1,{1\over 2}[f-\gamma^{2}r^{2}],\gamma,0\right)$ we see
$R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}=-\frac{\dot{f}}{r}+\gamma^{2}\left(1-f+{r^{2}f^{\prime\prime}\over 2}\right)=\frac{2\dot{m}}{r^{2}}+\gamma^{2}\left(2m^{\prime}-rm^{\prime\prime}\right)\,.$ (9)
The NCC requires that for all $\gamma$ we have $R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}\geq 0$, whence for $R(r,v)=r$ we deduce
$\mathrm{NCC}:\quad\dot{m}\geq 0;\qquad\mathrm{and}\qquad\left(2m^{\prime}-rm^{\prime\prime}\right)\geq 0\,.$ (10)
If (without prejudice as to any specific choice of dynamics) one chooses to define a function $\rho(r,v)$ by $m^{\prime}(r,v)=4\pi\rho(r,v)r^{2}$, then one sees $\left(2m^{\prime}-rm^{\prime\prime}\right)=-4\pi\rho^{\prime}(r,v)\,r^{3}$ and so in the current context the NCC in a generic null direction can be more compactly written as
$\mathrm{NCC}:\quad\dot{m}\geq 0;\qquad\mathrm{and}\qquad\rho^{\prime}\leq 0\,.$ (11)
II.2 Spacetimes with $R(r,v)$ having non-zero local minimum at $r=r_{0}$
If $\evaluated{R(r,v)}_{v=\text{const}}$ as a function of $r$ has a strict non-zero local minimum at some $r=r_{0}$, then the assumption that $R$ is non-negative implies $R_{0}\equiv R({r_{0}})>0$, in addition to both $R_{0}^{\prime}=0$ and $R_{0}^{\prime\prime}>0$. (Geometrically, this corresponds to a wormhole throat.) Consequently the expression in (4) evaluated at $r_{0}$ simplifies to
$\evaluated{R_{\mu\nu}\;k^{\mu}k^{\nu}}_{r=r_{0}}=-\frac{2R_{0}^{\prime\prime}}{R_{0}}\,.$ (12)
Independent of the kinematics, the NCC associated with the null vector $k^{\mu}$ tangent to ingoing null geodesics is always violated at least locally near $r=r_{0}$. [^5] More rarely, one might instead wish to consider $R(r,v)$ having a non-zero local maximum at $r=r_{0}$. Geometrically this would correspond to an anti-throat, a local maximum in the cross-sectional area of a “bag of gold” spacetime Fu:2019oyc ; Marolf:2008tx ; N-O-Murchadha_1987 . The NCC would be satisfied at such a local maximum. In a static spacetime where moreover $\dot{R}=0$ holds, the NCC associated with the null vector $l^{\mu}$ tangent to outgoing null geodesics is also violated locally near $r=r_{0}$, as can be seen from equation (5): $R_{\mu\nu}\;l^{\mu}l^{\nu}\to-f^{2}R^{\prime\prime}_{0}/(2R_{0})$. These observations apply in particular to static bouncing or wormhole geometries. A wormhole throat is defined such that the cross-sectional area of a bundle of light rays passing orthogonally through it has a strict local non-zero minimum. Therefore, if the throat is located at $r=r_{0}$, the function $R$ will satisfy the assumptions of this subsection. Static geometries with these properties in this sense by definition violate the NCC. This leads to a defocusing of both ingoing and outgoing null geodesics and ensures that the assumptions of the Penrose singularity theorem Penrose:1964wq are violated. Thereby bouncing or wormhole geometries can be free from spacetime singularities. We will explore the NCC violation expressed by equation (12) in the context of kinematical models which interpolate between a black hole with $R(r,v)=r$ and a wormhole or bouncing geometry.
Still in a static spacetime, in a completely general null direction, $\ell=\left(1,{1\over 2}[f-\gamma^{2}R^{2}],\gamma,0\right)$, at the throat one has
$R_{\mu\nu}\;\ell^{\mu}\ell^{\nu}\to-\left(f_{0}^{2}+{\gamma^{4}R_{0}^{4}}\right)\frac{R_{0}^{\prime\prime}}{2R_{0}}+{\gamma^{2}\over 2}\left(R_{0}^{2}f_{0}^{\prime\prime}+2\right).$ (13)
For both $\gamma\to 0$ and $\gamma\to\infty$ (the two radial null directions) we see that it is the sign of $R_{0}^{\prime\prime}$ that controls whether or not the NCC is violated at the throat/anti-throat.
III Kinematical transition models
In this section we will investigate possible violations of the NCC in kinematical processes describing transitions from a singular to a regular black hole or from a regular black hole to a horizonless compact object satisfying the assumptions of Section II.1. Subsequently we will demonstrate the NCC violation in transitions from a singular black hole satisfying the assumptions of Section II.1 to a regular bouncing or wormhole geometry satisfying the assumptions of Section II.2. To that end we first need to specify how we interpolate in time between two static geometries.
III.1 Interpolating mass and radial functions
Consider an initial static geometry $(i)$ at time $v_{i}$ and a final static geometry $(f)$ at time $v_{f}$ described by mass functions $m_{i,f}(r)$ and radial functions $R_{i,f}(r)$, possibly with implicit dependencies on some internal scale parameters $l_{i,f}$. Our goal is to define functions $m(r,v)$ and $R(r,v)$ interpolating in time between the two geometries. To that end we introduce an infinitely differentiable smooth transition function
$\sigma(v)=\frac{\tau(v)}{\tau(v)+\tau(1-v)}\,\,\,\quad\text{where}\quad\,\,\,\tau(v)=\begin{cases}e^{-\frac{1}{v}}&\text{for }v>0\,,\\ 0&\text{for }v\leq 0\,.\end{cases}$ (14)
This function satisfies $\sigma(v)=0$ for $v\leq 0$ and $\sigma(v)=1$ for $v\geq 1$ and $\dot{\sigma}(v)>0$ for $v\in(0,1)$, as shown in the left plot of Fig. 1. (While this specific choice of interpolating function is needed to quantitatively produce the figures below, it should be clear that any monotonic interpolating function with the same asymptotics will lead to qualitatively similar results.)
Figure 1: Left: Transition function $\sigma(v)$ (blue) defined in (14) and its time derivative $\dot{\sigma}(v)$ (green dashed). Right: Interpolating mass function $m(r,v)$ (blue) defined in (16) and its time derivative $\dot{m}(r,v)$ (green dashed) for an example with $m_{i}(r)<m_{f}(r)=2m_{i}(r)$.
Interpolating mass function: The interpolating mass function $m(r,v)$ should satisfy
$\lim_{v\to v_{i,f}}m(r,v)=m_{i,f}(r)\,\,\,\quad\text{and}\quad\,\,\,\lim_{r\to\infty}\dot{m}(r,v)=0\,.$ (15)
The last condition can be understood as demanding that the total ADM mass measured at infinity is preserved during the transition. Therefore, if $m_{i,f}(r)\to M_{i,f}$ for $r\to\infty$, we demand that $M_{i}=M_{f}\equiv M$. We can now use the function $\sigma(v)$ in (14) to switch between the mass function $m_{i}(r)$ at time $v_{i}\equiv 0$ and the mass function $m_{f}(r)$ at time $v_{f}\equiv 1$ by defining
$m(r,v)=\quantity(1-\sigma(v))m_{i}(r)+\sigma(v)m_{f}(r)\,,$ (16)
as illustrated in the right plot of Fig. 1.
Note that the time derivative of this mass function is
$\dot{m}(r,v)=\quantity(m_{f}(r)-m_{i}(r))\dot{\sigma}(v)\equiv\Delta m_{i\to f}(r)\dot{\sigma}(v)\,.$ (17)
Therefore $\dot{m}(r,v)\to\quantity(M_{i}-M_{f})\dot{\sigma}(v)=0$ for $r\to\infty$ and hence the last condition in (15) is satisfied. Since $\dot{\sigma}>0$ during the transition, for geometries satisfying the assumptions of Section II.1, i.e., for which $R(r,v)=r$, the NCC according to equation (8) is violated if
$\Delta m_{i\to f}(r)<0\,\,\,\Leftrightarrow\,\,\,m_{f}(r)<m_{i}(r)\,.$ (18)
We will provide examples for this NCC violation in transitions from singular black holes to regular black holes in Section III.2.
Now, if the functional form of the mass profile $m(r)$ for the initial and final geometries is identical up to a length parameter $l_{i,f}$, we may alternatively define an interpolating mass function $m(r,v)$ by making use of an interpolating length function $l^{n}(v)$,
$l^{n}(v)=\quantity(1-\sigma(v))l^{n}_{i}+\sigma(v)l^{n}_{f}\,,$ (19)
where $n$ is the power with which the parameter $l_{i,f}$ appears in the definition of the mass function $m(r)$ for the static initial and final geometry (e.g. later $n=2$ for Bardeen and $n=3$ for Dymnikova). The function $l^{n}(v)$ is then taken to be part of what defines the interpolating mass function by
$m(r,v)=m\quantity(r;l^{n}=l^{n}(v))\,.$ (20)
Note that by the chain rule the time derivative of this mass function is
$\dot{m}(r,v)=\quantity(\partial_{l^{n}}m)\partial_{v}\quantity(l^{n}(v))=\quantity(\partial_{l^{n}}m)\quantity(l_{f}^{n}-l_{i}^{n})\dot{\sigma}(v)\equiv\quantity(\partial_{l^{n}}m)\dot{\sigma}(v)\Delta l^{n}_{i\to f}\,.$ (21)
Therefore since $\dot{\sigma}>0$, and provided $\quantity(\partial_{l^{n}}m)<0$, for geometries satisfying the assumptions of Section II.1, i.e., for which $R(r,v)=r$, the NCC according to equation (8) is violated if
$\Delta l^{n}_{i\to f}>0\,\,\,\Longleftrightarrow\,\,\,l^{n}_{f}>l^{n}_{i}\,.$ (22)
We will provide examples for this NCC violation in transitions from regular black holes to horizonless compact objects in Section III.3.
Interpolating radial function: We can repeat the above constructions to define an interpolating radial function $R(r,v)$ such that
$\lim_{v\to v_{i,f}}R(r,v)=R_{i,f}(r)\,\,\,\quad\text{and}\quad\,\,\,\lim_{r\to\infty}R(r,v)\simeq r\,.$ (23)
The last condition imposes that at infinity the angular part of the line element (1) reduces to the usual area element for a 2-sphere of radius $r$. When combined with the last condition in (15) we therefore demand that at large distances the metric before, during and after the transition is well approximated by the asymptotically flat Schwarzschild metric with mass $M$. Using again the function $\sigma(v)$ in (14) to switch between the radial function $R_{i}(r)$ at time $v_{i}\equiv 0$ and the radial function $R_{f}(r)$ at time $v_{f}\equiv 1$, we define
$R(r,v)=\quantity(1-\sigma(v))\;R_{i}(r)+\sigma(v)\;R_{f}(r)\,.$ (24)
Note that the second radial derivative of this function is
$R^{\prime\prime}(r,v)=\quantity(1-\sigma(v))\;R_{i}^{\prime\prime}(r)+\sigma(v)\;R_{f}^{\prime\prime}(r)\,.$ (25)
Now if the initial geometry satisfies the assumptions of Section II.1, i.e., $R_{i}(r)=r$, then $R^{\prime\prime}_{i}=0$ and therefore, using that $\dot{\sigma}>0$, the NCC according to equation (4) will be violated if $R_{f}^{\prime\prime}>0$. We will provide examples for this NCC violation in transitions from singular black holes to bouncing or wormhole geometries in Section III.4.
Finally, if the functional form of the radial function $R(r)$ for the initial and final geometries is identical up a length parameter $l_{i,f}$, we may alternatively define an interpolating length function $R(r,v)$ by making use of the interpolating length function $l^{n}(v)$ in (19). Hence in this case we will define
$R(r,v)=R\quantity(r;l^{n}=l^{n}(v))\,.$ (26)
Note that the second radial derivative of this function is simply
$R^{\prime\prime}(r,v)=R^{\prime\prime}(r,l^{n}(v))\,.$ (27)
Therefore, if the initial and final geometries have $R^{\prime\prime}(r)>0$ the NCC will be be violated according to equation (4), independent of the value of the length parameters. We will provide an example for this NCC violation in a transition from a black bounce to a naked wormhole in Section III.5.
III.2 Transitions from singular to regular black holes
In the following subsubsections we illustrate the violation of the NCC in transitions from a singular to a regular black hole with $R(r,v)=r$ according to equation (8) using the interpolating mass function (16). The quantities of interest are the metric function $f(r,v)$ defined in (1) and the left hand side of the NCC in equation (8),
$f(r,v)=1-\frac{2m(r,v)}{r}\,\,\,\quad\text{and}\quad\,\,\,R_{\mu\nu}\;l^{\mu}l^{\nu}=\frac{2\dot{m}(r,v)}{r^{2}}\,.$ (28)
As the initial geometry we first choose a black hole with an integrable singularity for which the metric function $f(r,v)$ is finite at $r=0$ due to a positive length parameter $l_{i}$. We will also consider the limit $l_{i}\to 0$ corresponding to a Schwarzschild initial geometry. As the final geometries we choose two types of regular black holes with smooth de Sitter cores and length parameters denoted by $l_{f}$.
III.2.1 BH with integrable or Schwarzschild singularity $\longrightarrow$ Bardeen RBH
Fig. 2 shows the metric function $f(r,v)$ and left hand side of the NCC in equation (28) for an initial geometry corresponding to a black hole with an integrable singularity and a final geometry corresponding to a regular Bardeen black hole Bardeen:1968bh . The mass functions are given by
$m_{i}(r)=M\quantity(\frac{r}{r+l_{i}})\,\,\,\quad\text{and}\quad\,\,\,m_{f}(r)=M\quantity(\frac{r^{3}}{\quantity(r^{2}+l_{f}^{2})^{\frac{3}{2}}})\,.$ (29)
It is easy to check that $\left(2m^{\prime}-rm^{\prime\prime}\right)>0$ (that is, $\rho^{\prime}<0$) for these specific mass profiles, so (as per our previous discussion) for both the initial and final spacetimes the NCC is satisfied in all null directions, not just the radial null directions.
Fig. 3 shows the same quantities for $l_{i}\to 0$ when the initial geometry is Schwarzschild.
Figure 2: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a black hole with integrable singularity to a Bardeen regular black hole with mass functions (29) via the interpolating function $m(r,v)$ defined in (16). Along the red dashed lines $R_{\mu\nu}\;l^{\mu}l^{\nu}=0$, marking the edge of the NCC-violating region. Note that at any constant-$v$ slice the intersection with the vertical red dashed line is defined by the root of $\Delta m_{i\to f}(r)=0$ and therefore of $m_{i}(r)=m_{f}(r)$, according to equation (18). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. The length parameters are set to $l_{i}/M=1$ and $l_{f}/M=1/2$.
Figure 3: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a Schwarzschild black hole to a Bardeen regular black hole with mass functions (29) via the interpolating function $m(r,v)$ defined in (16). Along the red dashed horizontal lines $R_{\mu\nu}l^{\mu}l^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated in accordance with equation (18), as $m_{f}(r)<m_{i}(r)$ at any intermediate $v$. The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. The length parameters are set to $l_{i}/M=0$ and $l_{f}/M=1/2$.
III.2.2 BH with integrable or Schwarzschild singularity $\longrightarrow$ Dymnikova RBH
Fig. 4 shows the metric function $f(r,v)$ and left hand side of the NCC in equation (28) for an initial geometry corresponding to a black hole with an integrable singularity and a final geometry corresponding to a regular Dymnikova black hole Dymnikova:1992ux . The mass functions are given by
$m_{i}(r)=M\quantity(\frac{r}{r+l_{i}})\,\,\,\quad\text{and}\quad\,\,\,m_{f}(r)=M\quantity(1-e^{-\frac{r^{3}}{l_{f}^{3}}})\,.$ (30)
It is easy to check that $\left(2m^{\prime}-rm^{\prime\prime}\right)>0$ (that is, $\rho^{\prime}<0$) for these specific mass profiles, so (as per our previous discussion) for both the initial and final spacetimes the NCC is satisfied in all null directions, not just the radial null directions. Fig. 5 instead shows the same quantities for $l_{i}\to 0$ when the initial geometry is Schwarzschild.
The same qualitative features for the transition to a Dymnikova black hole can be observed as for the transition to a Bardeen black hole, which shows the insensitivity to the particular regular black hole final state.
Figure 4: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a black hole with integrable singularity to a Dymnikova regular black hole with mass functions (30) via the interpolating function $m(r,v)$ defined in (16). Along the red dashed lines $R_{\mu\nu}\;l^{\mu}l^{\nu}=0$, marking the edge of the NCC-violating region. Note that at any constant-$v$ slice the intersection with the vertical red dashed line is defined by the root of $\Delta m_{i\to f}(r)=0$ and therefore of $m_{i}(r)=m_{f}(r)$, according to equation (18). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. The length parameters are set to $l_{i}/M=1$ and $l_{f}/M=1$.
Figure 5: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a Schwarzschild black hole to a Dymnikova regular black hole with mass functions (30) via the interpolating function $m(r,v)$ defined in (16). Along the horizontal red dashed lines $R_{\mu\nu}l^{\mu}l^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated in accordance with equation (18), as $m_{f}(r)<m_{i}(r)$ at any intermediate $v$. The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. The length parameters are set to $l_{i}/M=0$ and $l_{f}/M=1$.
III.3 Transitions from regular black holes to horizonless compact objects
In the following we provide examples for the violation of the NCC according to equation (8) in transitions from an initial geometry describing a regular black hole with length parameter $l_{i}$ to a final geometry of the same type but with different value of this length parameter such that the spacetime describes a horizonless compact object. We denote the final value of the length parameter by $l_{f}$. To that end we now make use of the mass function (20) which uses the interpolating length function (19).
III.3.1 Bardeen RBH $\longrightarrow$ Bardeen HCO
For a transition from a regular black hole to a horizonless compact object based on a Bardeen geometry, we set the mass function to be
$m(r,v)=M\quantity(\frac{r^{3}}{\quantity(r^{2}+l^{2}(v))^{\frac{3}{2}}})\,.$ (31)
Fig. 6 shows the metric function $f(r,v)$ and left hand side of the NCC in equation (28) for an initial geometry corresponding to a Bardeen regular black hole and a final geometry corresponding to a Bardeen horizonless compact object. Note that both initial and final spacetimes satisfy the NCC, while the interpolating spacetime violates the NCC.
Figure 6: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a Bardeen regular black hole to a Bardeen horizonless compact object with interpolating mass function (31). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. Along the red dashed horizontal lines $R_{\mu\nu}l^{\mu}l^{\nu}=0$, and the NCC is satisfied. Everywhere else in between the initial and final geometries, the NCC is violated, in accordance with equations (21) and (22), as $\quantity(\partial_{l^{2}}m)<0$ and $l_{f}>l_{i}$. The length parameters are set to $l_{i}/M=1/2$ and $l_{f}/M=1$.
III.3.2 Dymnikova RBH $\longrightarrow$ Dymnikova HCO
For a transition from a regular black hole to a horizonless compact object based on a Dymnikova geometry, we set the mass function to be
$m(r,v)=M\quantity(1-e^{-\frac{r^{3}}{l_{f}^{3}(v)}})\,.$ (32)
Fig. 7 shows the metric function $f(r,v)$ and the NCC in equation (28) for an initial geometry corresponding to a Dymnikova regular black hole and a final geometry corresponding to a Dymnikova horizonless compact object. Note that both initial and final spacetimes satisfy the NCC, while the interpolating spacetime violates the NCC. As before we observe the same qualitative features for the transition to a Dymnikova HCO as for the transition to a Bardeen HCO, which shows the insensitivity to the particular HCO final state.
Figure 7: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (28) for a transition from a Dymnikova regular black hole to a Dymnikova horizonless compact object with interpolating mass function (32). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. Along the red dashed horizontal lines $R_{\mu\nu}l^{\mu}l^{\nu}=0$, and the NCC is satisfied. Everywhere else in between the initial and final geometries, the NCC is violated, in accordance with equations (21) and (22), as $\quantity(\partial_{l^{3}}m)<0$ and $l_{f}>l_{i}$. The length parameters are set to $l_{i}/M=1/2$ and $l_{f}/M=3/2$.
III.4 Transitions from singular black holes to bouncing or wormhole geometries
In the following we provide examples for the violation of the NCC in transitions from a black hole with an integrable singularity satisfying the assumptions of Section II.1, to a regular Simpson–Visser geometry Simpson:2018tsi ; Lobo:2020ffi ; Franzin:2021vnj ; Lobo:2020kxn satisfying the assumptions of Section II.2, according to equation (4). We use the interpolating mass function $m(r,v)$ in (16) and radial function $R(r,v)$ in (24), and set
$m_{i}(r)=M\quantity(\frac{r}{r+l_{i}})\,\,\,\quad\text{and}\quad\,\,\,M_{f}(r)=M\,,$ (33)
as well as
$R_{i}(r)=r\,\,\,\quad\text{and}\quad\,\,\,R_{f}(r)=\sqrt{r^{2}+l_{f}^{2}}\,.$ (34)
The quantities of interest are the metric function $f(r,v)$ defined in (1) and the left hand side of the NCC in equation (4),
$f(r,v)=1-\frac{2m(r,v)}{R(r,v)}\,\,\,\quad\text{and}\quad\,\,\,R_{\mu\nu}k^{\mu}k^{\nu}=-\frac{2R^{\prime\prime}(r,v)}{R(r,v)}\,.$ (35)
Note that now the initial spacetime satisfies the NCC, while the final spacetime violates the NCC. So it is now not too surprising that the interpolating spacetime also violates the NCC.
III.4.1 BH with integrable or Schwarzschild singularity $\longrightarrow$ SV black bounce
Fig. 8 shows the expressions (35) for an initial geometry with parameter $l_{i}$ such that the spacetime contains a black hole with an integrable singularity and a final parameter $l_{f}$ such that the spacetime contains a Simpson–Visser black bounce. Fig. 9 shows the same quantities for $l_{i}\to 0$ when the initial geometry is Schwarzschild.
Figure 8: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (35) for a transition from a black hole with integrable singularity to a Simpson–Visser black hole with mass and radial functions (33) and (34), via the interpolating mass and radial functions (16) and (24). The gray-shaded region marks the trapped region where $f(r,v)<0$. Along the red dashed horizontal line $R_{\mu\nu}k^{\mu}k^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated, in accordance with equation (27). The length parameters are set to $l_{i}/M=1$ and $l_{f}/M=1$.
Figure 9: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (35) for a transition from a Schwarzschild black hole to a Simpson–Visser black hole with mass and radial functions (33) and (34), via the interpolating mass and radial functions (16) and (24). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the left plot is bounded by the minimum value of the legend and represents a region where the lapse function $f(r,v)$ becomes large negative and off-scale.Along the red dashed horizontal line $R_{\mu\nu}k^{\mu}k^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated, in accordance with equation (27). The length parameters are set to $l_{i}/M=0$ and $l_{f}/M=1$.
III.4.2 BH with integrable or Schwarzschild singularity $\longrightarrow$ SV naked wormhole
Fig. 10 shows the expressions (35) for an initial geometry with parameter $l_{i}$ such that the spacetime contains a black hole with an integrable singularity and a final parameter $l_{f}$ such that the spacetime contains a Simpson–Visser naked wormhole. Fig. 11 shows the same quantities for $l_{i}\to 0$ when the initial geometry is Schwarzschild.
Figure 10: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (35) for a transition from a black hole with integrable singularity to a Simpson–Visser naked wormhole with mass and radial functions (33) and (34), via the interpolating mass and radial functions (16) and (24). The uncolored region in the left plot is bounded by the minimum value of the legend and represents a region where the lapse function $f(r,v)$ becomes large negative and off-scale. The gray-shaded region marks the trapped region where $f(r,v)<0$. Along the red dashed horizontal line $R_{\mu\nu}k^{\mu}k^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated, in accordance with equation (27). The length parameters are set to $l_{i}/M=1$ and $l_{f}/M=5/2$.
Figure 11: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (35) for a transition from a Schwarzschild black hole to a Simpson–Visser black hole with mass and radial functions (33) and (34), via the interpolating mass and radial functions (16) and (24). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the left plot is bounded by the minimum value of the legend and represents a region where the lapse function $f(r,v)$ becomes large negative and off-scale. Along the red dashed horizontal line $R_{\mu\nu}k^{\mu}k^{\nu}=0$, and the NCC is satisfied. Everywhere else the NCC is violated, in accordance with equation (27). The length parameters are set to $l_{i}/M=0$ and $l_{f}/M=5/2$.
III.5 Transitions from bouncing geometries to wormhole geometries
In the following, we provide examples for the violation of the NCC according to equation (4) from an initial geometry describing a regular bouncing geometry with length parameter $l_{i}$ to a final geometry of the same type but with different values of this length parameter such that the spacetime describes a regular naked wormhole, whose length parameter we denote by $l_{f}$. To that end we use the interpolating mass function defined as in equation (16) together with the interpolating radial function (26). For concreteness, we focus on a transition model based on a Simpson–Visser geometry Simpson:2018tsi ; Lobo:2020ffi ; Franzin:2021vnj ; Lobo:2020kxn .
III.5.1 SV black bounce $\longrightarrow$ SV naked wormhole
For the transition from a Simpson–Visser black-bounce spacetime to a Simpson–Visser wormhole spacetime Simpson:2018tsi ; Lobo:2020ffi ; Franzin:2021vnj ; Lobo:2020kxn we use the interpolating mass function $m(r,v)$ in (16) and set
$m_{i}(r)=m_{f}(r)=M\,.$ (36)
Moreover, according to equation (26) we set the radial function $R(r,v)$ to be
$R(r,v)=\sqrt{r^{2}+l^{2}(v)}\,,$ (37)
where $l^{2}(v)$ is defined in (19). As before, the quantities of interest are the metric function $f(r,v)$ defined in (1) and the left hand side of the NCC in equation (4) as summarized in equation (35). Fig. 12 shows these expressions for an initial geometry with parameter $l_{i}$ such that the spacetime contains a Simpson–Visser black hole and a final parameter $l_{f}$ such that the spacetime contains a Simpson–Visser naked wormhole. Note that now both the initial and final spacetimes violate the NCC. So it is now not too surprising that the interpolating spacetime also violates the NCC.
Figure 12: Plots of the metric function $f(r,v)$ (left) and the NCC (right) in (35) for a transition from a Simpson–Visser black hole to a Simpson–Visser naked wormhole with mass functions (36) via the interpolating mass and radial functions (16) and (37). The gray-shaded region marks the trapped region where $f(r,v)<0$. The uncolored region in the right plot is bounded by the minimum value of the legend and represents a region of extremely high and off-scale NCC violation. The length parameters are set to $l_{i}/M=1$ and $l_{f}/M=5/2$.
IV Discussion
So how do regular black holes evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems?
As pure mathematics, one can certainly write down many line elements that describe regular black holes all satisfying the NCC globally. This is certainly the case for the standard regular black hole spacetimes, [^6] If desired, these regular black hole “models” can sometimes be bootstrapped to “solutions” by reverse engineering a suitable matter Lagrangian — typically some variant on nonlinear electrodynamics. But this is not the main focus of the current article. such as Bardeen, Dymnikova, Hayward, and their close variants Bardeen:1968bh ; Dymnikova:1992ux ; Hayward:2005gi . [^7] More distantly related regular black holes, such as the black-bounce solutions (or the closely related naked traversable wormholes) do however violate the NCC globally. The intermediate “Minkowski core” regular black holes satisfy the NCC in the deep core, but violate it in the asymptotic region.
So, how do regular black holes actually form? One possibility is simply that the Hawking–Penrose singularity theorems cannot be applied to static or stationary regular black-hole gemoetries, as they are endowed, (in the same way as the singular Reissner–Nordström, Kerr, and Kerr–Newman spacetimes), with inner horizons which are also Cauchy horizons, and as such violate global hyperbolicity. This would, however, be a quite debatable resolution of the conundrum, as such geometries have to be considered as more mathematical abstractions, asymptotic states, rather than realistic models of the end points of a gravitational collapse.
(To put this in a somewhat blunter manner: The standard Hawking–Penrose singularity theorems are of limited utility in the presence of static or stationary inner horizons, including the Reissner–Nordström, Kerr, and Kerr–Newman spacetimes Senovilla:2014gza . Yes, there are finite affine-length null geodesics, but they accumulate on the Cauchy horizon — not on the singularity.)
On the other hand, in an evolving time-dependent situation, an inner horizon does not necessarily need to be a Cauchy horizon, and so it could be compatible with the assumption of global hyperbolicity. In such a situation, it seems one would be forced to violate the NCC in order to reach a singularity-free end point.
We have therefore developed a number of kinematic models that interpolate between “standard physics” (standard black holes or horizonless objects) and regular black holes (assuming spherical symmetry), and have explicitly verified that during the purely kinematic evolution process (no particular dynamics is assumed) violations of the NCC are quite common and apparently unavoidable, although a formal proof of the universality of this behaviour would certainly be desirable.
We feel that this investigation serves to highlight more than ever the pressing need for improved singularity theorems, better-adapted to dealing with evolving inner and outer trapping horizons Hawking:2014tga ; Visser:2014zqa ; Ashtekar:2004cn . In particular, our investigations emphasize the need for theorems which are able to provide general conditions under which the formation of timelike singularities — shielded by dynamical inner horizons — is unavoidable. In conclusion, one might say that also in this situation, as has recently happened for mass inflation Carballo-Rubio:2024dca , regular black holes are forcing us to carefully reconsider the dynamical aspects of trapped regions and their horizons. As such, independently of their phenomenological interest, they ought to be considered crucial objects for current and future research in black hole physics (see e.g. Carballo-Rubio:2023mvr ; DiFilippo:2024spj ; Buoninfante:2024oxl ; Visser:2024zkx ; Carballo-Rubio:2025fnc ).
Acknowledgements.
JB is supported by a doctoral scholarship by the German Academic Scholarship Foundation (Studienstiftung). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada, and by the Province of Ontario through the Ministry of Colleges and Universities.
During early phases of this work MV was supported by a Victoria University travel grant, and by SISSA and INFN.